*10. Möbius transformations. Let a,b,c,d EC with ad-bc70. The function T(2) = 2 a2 + b cz + d à (2 +-d/c) is called a Möbius10. Mobius transformations. Let a, b, c, d \in ad-bc \neq 0 . The function is called a Mobius transformation (or linear fractional transformation). Show that

a) lim z->inf T(z) = inf if c=0;

b)kim z-> inf T(z) = a/c and lim z-> d/c T(z) = inf if c\neq0
Answer #1

Mobius transformation

T(z) = 9z+b z A - Fud/c. Cz+ 17 If C=0. T(z) = az +b = 22 +b lim T(z) = lim 9z+b - OD Z- 2 -) . ] lim 270 (z) = lim ztb 27 c